### v1

# TrackPGD: A White-box Attack using Binary Masks against Robust Transformer Trackers

Abstract
Object trackers with transformer backbones have achieved robust performance on visual object tracking datasets. However, the adversarial robustness of these trackers has not been well studied in the literature. Due to the backbone differences, the adversarial white-box attacks proposed for object tracking are not transferable to all types of trackers. For instance, transformer trackers such as MixFormerM still function well after black-box attacks, especially in predicting the object binary masks. We are proposing a novel white-box attack named TrackPGD, which relies on the predicted object binary mask to attack the robust transformer trackers. That new attack focuses on annotation masks by adapting the well-known SegPGD segmentation attack, allowing to successfully conduct the white-box attack on trackers relying on transformer backbones. The experimental results indicate that the TrackPGD is able to effectively attack transformer-based trackers such as MixFormerM, OSTrackSTS, and TransT-SEG on several tracking datasets.
1 Introduction
Adversarial attacks [1, 2, 3, 4] can significantly harm neural networks performances by adding carefully crafted imperceptible noise to the input. For instance, classification networks [5, 6] are highly susceptible to such adversarial perturbations. For object segmentation and object tracking, some recent attacks [7, 1, 2, 4, 3] have been developed to disrupt the object masks and bounding boxes output, respectively. Since 2020, the Visual Object Tracking (VOT) challenge [8] introduced a new sub-challenge that requires object trackers to present binary masks as output and be evaluated based on their mask generation ability. In addition, novel trackers with transformer backbones [9, 10, 11] have emerged, whose approaches infer the object mask, bounding box, and confidence score as a triple output. As a result, object binary masks have become increasingly important both as a valid output of trackers and a valid criterion for evaluation in tracking benchmarks.
Visual tracking can be compromised by adversarial attacks in different ways, by manipulating elements such as object bounding boxes [2], classification labels [3, 4, 1], and regression labels [3, 1]. Like other adversarial attacks, the object tracker’s attacks can be categorized into two categories: white-box attacks [7, 3], where the attacker has access to the model’s gradients, and black-box attacks [2], where only the tracker’s output is available. In Table 1, several attack methods and their attack proxy are listed along with their applicability to some new transformer tracker, MixFormerM [9], OSTrackSTS [10] and TransT-SEG [11].
Method Attack Proxy MixFormerM [9] OSTrackSTS [10] TransT-SEG [11]
BB IoU [2] Object bbox $\checkmark$ $\checkmark$ $\checkmark$
CSA [4] Object bbox, heat-maps $\checkmark$ $\checkmark$ $\checkmark$
WB SPARK [1] Regr. and classif. labels $\times$ $\times$ $\checkmark$
RTAA [3] Regr. and classif. labels $\times$ $\times$ $\checkmark$
TrackPGD Object binary mask $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Characterization of attacks methods on transformer-based trackers. TrackPGD is the only applicable White-Box (WB) attack against MixFormerM, and OSTrackSTS, using a new proxy (i.e. binary mask), while the two other methods that can be applied to these trackers, IoU and CSA, are Black-Box (BB) methods manipulating object bounding box.
Figure 1: The binary masks generated by MixFormerM [9] and OSTrackSTS [10] attacked by the IoU [2] vs. TrackPGD. The green/red masks are the before and after the attack outputs.
Table 1 presents several attacks along with their respective attack proxies and specifies their applicability to new transformer trackers. For instance, MixFormerM [9] does not output the labels of object candidates. Therefore, the attack algorithms with the objects labels proxy, such as RTAA [3] and SPARK [1], are not applicable to MixFormerM as a white-box attack. In this paper, we present a white-box attack, named TrackPGD, to attack the transformer trackers via a new proxy, that is the object binary mask. Figure 1 indicates TrackPGD performance on crashing binary masks generated by MixFormerM [9] and OSTrackSTS [10]. The green/red masks represents the predicted mask before and after the attack, respectively. As a white-box attack applicable on both MixFormerM and OSTrackSTS, the binary mask after the attack (red mask) caused more damage than the other applicable attack, i.e. Intersection over Union (IoU) method [2]. Although the IoU method may degrade the MixFormerM’s bounding box, the binary mask remained almost unchanged. Yet, the bounding box can be obtained from the object binary mask. Furthermore, a new tracking dataset such as VOT2022STS [12] evaluates the tracker’s performance based on only the generated mask, making the target object binary masks key elements to consider in order to decrease the robustness of the tracker. Our contributions can be summarized as follows:
  • We propose a novel white-box attack, TrackPGD, which builds the adversarial noise from the binary mask to attack transformer trackers.
  • A new loss term – built upon SegPGD attack – is proposed to compute the TrackPGD loss with the aim of misleading visual trackers in providing an accurate binary mask.
  • MixFormerM experienced significant accuracy ($-75\%$), average overlap ($-97\%$) and robustness ($-91\%$) reductions after applying TrackPGD, evaluated on the VOT2022STS dataset.
  • Experimental results also demonstrate that the perturbations generated by TrackPGD have a significant influence on bounding box predictions in tracking benchmarks.
2 Related works
Robust Object Trackers Tracker robustness was initially defined as the frequency of failures, marked by the overlap between prediction and ground truth dropping to zero. This definition is employed in datasets like VOT2016 [13] and VOT2018 [14]. In VOT evaluation for short-term tracking, the tracker is reset after each failure. A newer protocol introduced in VOT2020 [8] includes multiple anchors in each test sequence spaced approximately 50 frames apart to facilitate anchor-based short-term tracking [8, 12]. This protocol measures robustness based on the duration of video subsequences where the overlap falls below a fixed threshold for several frames. In contrast, other tracking datasets, like LaSOT benchmark [15], conduct a one-pass evaluation. The three most robust trackers on the VOT2022STS [12] were the Multi-Scale approach based on Object associating with Transformers (MS-AOT) [16, 17], Distractor-Aware Multi-head Tracker (DAMTMask) [18, 19], and MixFormerM [9]. MS-AOT used MixFormer [9] for targets smaller than $1/900$ frame size and for larger targets, a multi-scale object association approach is employed to perform tracking task. DAMTMask integrates MixFormer [9] and TransT-SEG [11] to construct a robust performance. The MixFormer [9] tracker proposed a Mixed Attention Module (MAM) to jointly extract features of search and template regions. MixFormerM [9, 12] contains two networks, one tracker network and the other is AlphaRefine [20] for object segmentation. But TransT-SEG [11] jointly predicts the object bounding box and generates the object binary mask via one single network that uses a light cross-attention mechanism to fuse deep features. Another transformer-based tracker, OSTrackSTS [10], gained the third rank on the VOT-STS2022 [12] sequestered dataset after MS-AOT and DAMTMask. The OSTrackSTS [10] is the One-Stream Track method [10] coupled with the AlphaRefine [20] to provide object binary masks.
Adversarial attacks for object trackers The attack proxy can be classification labels [3, 4, 1], regression output [3] or even the motion features of the object [2, 3]. In the Robust Tracking against Adversarial Attacks (RTAA) method [3], specific labels for classification and regression are modified. An online incremental attack called SPARK [1] has been developed for object trackers that use the previous frame perturbations and manipulate the object classification and regression labels to generate the adversarial frame. The Cooling-Shrinking Attack (CSA) method, as described in [4], involves training two Generative Adversarial Networks (GANs) to create adversarial images for the SiamRPN++ tracker [21]. The cooling term cools down the heated pixels in the heatmap of SiamRPN++, while the shrinking term shrinks the predicted bounding box. The IoU attack is a black-box approach that uses the predicted bounding box to inject two types of noises [2] into frames. The directional and orthogonal perturbations are injected in a given frame to degrade the predicted bounding box from the previous prediction.
3 Method
3.1 Background: SegPGD
SegPGD [7] loss for a segmentation network ($L_{\text{SegPGD}}$) is defined to steer the neural network in providing an inaccurate segmentation map. This loss is computed between the mask annotation $Y$ and the adversarial segmentation map $f_{\text{seg}}(X^{\text{adv}})$ generated by the segmentation network from the adversarial image $X^{\text{adv}}$:
$L_{\text{SegPGD}}(f_{\text{seg}}(X^{\text{adv}}),Y)=\frac{1-\lambda}{HW}L_{\text{pos}}+\frac{\lambda}{HW}L_{\text{neg}}\,,$ (1)
where $\lambda$ is a coefficient that dynamically changes over iterations and $H$ and $W$ are the frame’s height and width, respectively. Here, $L_{\text{pos}}$ and $L_{\text{neg}}$ denote the cross-entropy loss restricted to correctly classified and wrongly classified pixels, respectively. The motivation for this reweighting stems from the observation that gradients of cross-entropy for wrongly classified examples tend to have larger amplitude, while the attack aims primarily at correctly classified pixels [7].
To adapt SegPGD for use against object trackers and to target the predicted binary masks, we must overcome two key challenges. Firstly, while SegPGD operates on multi-class segmentation networks, binary masks only have two classes, necessitating the switching of classification labels between pixels. Secondly, the labels in the majority of samples are imbalanced, with the number of object pixels being significantly fewer than background pixels. To address these hurdles and effectively attack trackers, we introduce TrackPGD.
3.2 Proposed method: TrackPGD
Let us formalize object tracking as using a tracker $\mathcal{F}(\cdot)$ that processes a video sequence $\mathcal{V}=\{I_{0},\ldots,I_{L}\}$ with $L+1$ frames, having the first frame object mask $M_{0}$ provided to initialize the tracker. The transformer tracker is expected to estimate the object mask $M_{\tau}$ at each time step $\tau\in\{1,\ldots,L\}$ from the given video frame $I_{\tau}$. The goal of the attack method is thus to create an adversarial video frame $I^{\text{adv}}_{\tau}$ at each time step to mislead the object tracker in its mask prediction. The transformer tracker predicts at each time $\tau$ a mask $M_{\tau}$ which assigns to each pixel a probability of being an object. The TrackPGD attack aims to mislead the tracker in making these assignments by creating the adversarial frame $I^{\text{adv}}_{\tau}$.
For object tracking, the mask annotation $G_{\tau}$ consists of two labels (object (1) and background (0)) and it follows that finding adversarial examples can be achieved at optimization step $t$ both by maximizing $L_{\text{SegPGD}}(M^{t},G_{\tau})$ or by minimizing $L_{\text{SegPGD}}(M^{t},1-G_{\tau})$, where we consider the mask $M_{\tau-1}$ predicted on the previous frame as the ground truth $G_{\tau}$ at step $\tau$. In the TrackPGD framework, we therefore decide to combine these two complementary objectives into a difference loss $L_{\Delta}$ as follows:
$L_{\Delta}=L_{\text{SegPGD}}(M^{t},G_{\tau})-L_{\text{SegPGD}}(M^{t},1-G_{\tau})\,.$ (2)
Since the object pixels and background pixels are imbalanced in the video frame, we employed the focal loss [22]. It is a useful loss for training highly imbalanced data, which is the case, here, for the object and background pixels. For TrackPGD, instead of using cross-entropy as in the original equation of focal loss [22], we use the $L_{\Delta}$ loss as follows:
$p_{t}$ $=G_{\tau}\mathcal{P}+(1-G_{\tau})(1-\mathcal{P})$
$L_{\text{focal}}$ $=\alpha_{t}(1-p_{t})^{\gamma}L_{\Delta}\,,$ (3)
where $\mathcal{P}$ is a probability map derived from the application of the sigmoid function on the predicted binary mask $M_{t}$(before thresholding) and $\gamma$ is the focusing parameter. This $\alpha$-balanced variant of the focal loss used is similar to the TransT-SEG [11] except that $L_{\Delta}$ is used instead of cross-entropy. Then, we compute the dice loss [23] between $M^{t}$ and $G_{\tau}$ as follows:
$L_{\text{dice}}=1-2\text{IoU}(M^{t},G_{\tau}),$ (4)
where IoU is the Intersection over Union loss between the ground truth $G_{\tau}$ and the predicted binary mask $M^{t}$. The combination of focal loss (Section 3.2), and dice loss (Eq. 4), is used to compute the TrackPGD loss as follows:
$L=\lambda_{1}L_{\text{focal}}+\lambda_{2}L_{\text{dice}}\,,$ (5)
where the $\lambda_{1}$ and $\lambda_{2}$ balance the focal and dice terms. The pseudocode of TrackPGD is presented in Algorithm 1. The sign of step size $\alpha$, line 9 in Algorithm 1, is determined according the best performance of each tracker, following the decoupling of the norm and direction of gradients in adversarial attacks [24].
Algorithm 1 TrackPGD to attack transformer trackers with segmentation capability
1:Tracker $\mathcal{F}(\cdot)$, current frame ${I}_{\tau}$, previous binary mask ${M}_{\tau-1}$, perturbation range $\epsilon$, step size $\alpha$, loss trade-offs $\lambda_{1}$ and $\lambda_{2}$, maximum iteration ${T}$, focusing parameter $\gamma$, variant of focal loss $\alpha_{t}$, probability map $p_{t}$
2:${I}_{\text{adv}}^{0}\leftarrow I_{\tau}$triangleright initialization
3:$G_{\tau}\leftarrow M_{\tau-1}$triangleright use last predicted binary mask as ground truth
4:for $t=1\ldots T$ do
5:     ${M}^{t}\leftarrow\mathcal{F}({I}_{\text{adv}}^{t-1})$triangleright predict binary mask
6:     $L_{\Delta}\leftarrow L_{\text{SegPGD}}(M^{t},G_{\tau})-L_{\text{SegPGD}}(M^{t},1-G_{\tau})$
7:     $L_{\text{focal}}\leftarrow\alpha_{t}(1-p_{t})^{\gamma}L_{\Delta}$
8:     $L_{\text{dice}}\leftarrow 1-2\,\text{IoU}(M^{t},G_{\tau})$
9:     $L\leftarrow\lambda_{1}L_{\text{focal}}+\lambda_{2}L_{\text{dice}}$triangleright compute TrackPGD loss
10:     ${I}_{\text{adv}}^{t}\leftarrow I_{\text{adv}}^{t-1}+\alpha\,\mathrm{sign}(\nabla_{I_{\text{adv}}^{t-1}}L)$
11:     ${I}_{\text{adv}}^{t}\leftarrow\phi^{\epsilon}\left({I}_{\text{adv}}^{t}\right)$triangleright clip to the $\epsilon$-ball
12:end for
4 Experiments
We aim to examine TrackPGD’s impact on transformer tracker performance compared to other attack methods. We will start by outlining our experimental setup and protocols. Next, we will evaluate TrackPGD’s effectiveness on transformer trackers using object masks. The ablation study will illustrate the significance of our proposed difference loss for tracking and the impact of TrackPGD loss terms on overall performance. Additionally, we will showcase evaluation results based on predicted bounding boxes, highlighting TrackPGD’s advantages.
Evaluation setup
TrackPGD is compared with other existing attacks on VOT-STS2022 [12], VOT2018 [14] and VOT2016 [13] datasets. We conducted two primary experiments containing baseline, and unsupervised, on VOT-STS2022. In the baseline experiments, the Expected Overlap Analysis (EOA) and Accuracy-Robustness analysis are conducted using the anchor-based short-term tracking protocol. The unsupervised experiment computes the average accuracy in one initialization setting per sequence. The VOT2018 [14] and VOT2016 [13] are annotated with object bounding boxes and used here to evaluate the tracker’s robustness in bounding box prediction after applying the attacks. We selected three popular transformer-based trackers, MixFormerM [9], OSTrackSTS [10] and TransT-SEG [11], to assess their adversarial robustness against adversarial attacks. We applied four attack methods including CSA [4], SPARK [1], RTAA [3], and IoU [2], against these trackers. The TrackPGD perturbs the search region of the segmentation network in MixFormerM [9] and OSTrackSTS [10] while the search region of the tracking network remains untouched. The hyperparameters $(\lambda_{1},\lambda_{2})$ of the TrackPGD loss are determined through the validation experiment on the DAVIS2016 dataset [25], as explained in the Supplementary Materials.
4.1 Binary mask assessment
We evaluated the TrackPGD attack compared to other methods on two transformer trackers – MixFormerM [9] and OSTrackSTS [10] – using the VOT-STS2022 dataset, see  Table 2. Two streams of experiments from VOT-STS2022 (baseline and unsupervised) are conducted to indicate our attack performance under various protocols and metrics. Our findings demonstrate that TrackPGD was the most effective attack, reducing the robustness of MixFormerM [9] by around $91\%$ in the baseline experiment of VOT-STS2022 evaluation. TrackPGD also caused higher degradations to other evaluation metrics – EAO, accuracy, and AUC – than the two attack methods (i.e., IoU [2] and CSA [4] methods) against MixFormerM [9], both the on baseline and unsupervised streams of experiments. Among all measured metrics, TrackPGD ranked first by a significant margin on the VOT-STS2022 dataset. In attacking the OSTrackSTS tracker, TrackPGD also ranked first as an applicable attack per EAO, accuracy and AUC metrics. However, the IoU method [2] performs better than our method on the robustness metric. Figure 2 represents the performance of TrackPGD in generating inaccurate binary masks for MixFormerM [9] and OSTrackSTS [10]. The green color represents the original trackers’ output, while the red color represents the tracker output after the TrackPGD attack.
[…]
Attack EAO Acc. Robust. AUC EAO Acc. Robust. AUC
MixFormerM No Attack 0.59 0.80 0.88 0.72
CSA 0.56 0.80 0.86 0.65 4.58 -0.63 2.61 9.40
IoU 0.36 0.66 0.68 0.51 39.05 17.29 23.07 15.40
TrackPGD 0.02 0.20 0.08 0.06 96.94 75.44 91.14 91.70
OSTrackSTS No Attack 0.59 0.77 0.87 0.66 - - - -
CSA 0.53 0.76 0.82 0.62 10.54 2.32 5.29 5.79
IoU 0.31 0.71 0.60 0.48 47.45 8.13 30.92 27.94
TrackPGD 0.23 0.31 0.86 0.27 60.34 59.74 0.80 59.06
Table 2: The MixFormerM [9] and OSTrackSTS [10] results in generating object binary masks, when attacked by different methods on the VOT-STS2022 [12] dataset. Bold/italic are the first/second ranks, respectively.
Figure 2: The results of MixFormerM [9] and OSTrackSTS [10] attacked by TrackPGD. The green mask is the original output, while the red mask is the output after attack.
4.2 Ablation study
Role of $L_{\Delta}$ in TrackPGD
Inspired by the SegPGD attack [7], TrackPGD relies on the predicted binary mask to generate the adversarial frames. The main challenge in following SegPGD for binary mask impairment to attack object trackers is that the SegPGD works well as a multi-class segmentation attack; while here, we have only two classes to switch the pixel labels. The object class and the background class pixels should switch their labels after TrackPGD. However, the object pixels are limited in comparison to the background pixels in almost all of the video frames. For this specific reason, it is necessary to define our method based on both object and background masks to involve all of the frame pixels in generating the adversarial perturbation. Plus, given the availability of two classes in the tracking scenario, we customized the segmentation loss using the tracking loss of the TransT [26] to effectively attack binary masks. We also demonstrate the role of ${L}_{\Delta}$ in TrackPGD versus using the vanilla SegPGD losses for the object and background masks, in Figure 3. The after attack masks show the essential role of our proposed combination loss $L_{\Delta}$ in crashing the object mask inferred by MixFormerM [9]. Although, the vanilla SegPGD losses also generate inaccurate binary masks, the mask impairment caused by $L_{\Delta}$ is significantly greater.
$\quad L_{\text{SegPGD}}(M^{t},G_{\tau})$ $-L_{\text{SegPGD}}(M^{t},1-G_{\tau})$ $\qquad\quad L_{\Delta}$
Figure 3: The before (green) and after attack masks (red) generated by MixFormerM [9], while the vanilla SegPGDs and difference loss are used instead of binary cross entropy in the focal loss for the TrackPGD attack.
In addition, we conducted an ablation study to demonstrate the difference loss role in the focal loss. We predicted the binary masks by the MixFormerM tracker attacked by TrackPGD where the binary cross entropy is substituted with the vanilla SegPGD loss for the object mask, vanilla SegPGD for the background mask, and the difference loss $L_{\Delta}$. The DAVIS2016 [25] dataset was used to compute the percentage of the mean Jaccard index $J(M)$ of the binary masks indicating the average overlap between predicted and annotated binary masks. TrackPGD attack is stronger with the proposed difference loss, see Table 3, in comparison with the vanilla SegPGD losses for the object mask $G_{\tau}$, and the background mask $1-G_{\tau}$. The difference loss presented the most effective attack based on the generated binary masks by MixFormerM [9], which has the smaller $J(M)\%$ on average.
Original $L_{\text{SegPGD}}(M^{t},G_{\tau})$ $-L_{\text{SegPGD}}(M^{t},1-G_{\tau})$ $L_{\Delta}$
$J(M)\%$ 85.82 52.86 37.28 30.30
Table 3: Cross-entropy substitutions in focal loss. The vanilla SegPGD losses versus the difference loss in the TrackPGD against MixFormerM [9] on DAVIS2016 [27].
Coefficients of the TrackPGD Loss
In this study, we show the effect of each loss term, dice and focal losses, in the TrackPGD attack. We conducted experiments on DAVIS2016 [25] using two evaluation metrics: the mean Jaccard index $J(M)$ and contour accuracy $F(M)$ of the predicted masks. The results were presented in Table 4 where the average $J\&F(M)$ percentage was used to demonstrate the impact of each losses term, focal vs. dice loss. Table 4 shows that the focal loss is more effective than the dice loss, resulting in smaller $J\&F(M)$ values. The $J\&F(M)$ value is used to measure the effectiveness of the attack, where a smaller value represents a more significant influence. We selected powers of 10 as coefficients of focal loss and powers of 2 as coefficients of dice loss. As in the literature, the loss related to IoU is normally 1 or 2 [11, 9], while the loss terms related to binary cross entropy (or focal loss) are selected as 1000 [20] or 1 [11]. An interesting observation in  Table 4 is that the MixFormerM is significantly impacted by TrackPGD, more than the TransT-SEG tracker using the same set of experiments. In both tables, the MixFormerM tracker achieves the smallest values of $J\&F(M)$ meaning more significant attack results.
$\lambda_{1}$ (with $\lambda_{2}=0$) $10000$ $1000$ $100$ $10$ $1$ $0.1$
MixFormerM 27.97 27.87 27.76 27.90 27.73 27.93
TransT-SEG 28.98 28.52 28.74 28.71 28.65 28.67
$\lambda_{2}$ (with $\lambda_{1}=0$) $4$ $2$ $1$ $0.5$ $0.25$ $0.125$
MixFormerM 30.00 30.27 29.92 30.04 30.14 29.92
TransT-SEG 33.62 33.62 33.52 34.05 33.14 33.55
Table 4: $J\&F(M)\%$ values with varying focal and dice losses. Present the role of focal loss (changing $\lambda_{1}$ when $\lambda_{2}=0$) and dice loss (changing $\lambda_{2}$ when $\lambda_{1}=0$) in the TrackPGD attack. $J\&F(M)\%$ values are computed on DAVIS2016 [27] dataset.
4.3 Bounding box assessment
In  Table 5, we compare the performance of TrackPGD with other methods against the TransT-SEG [11] tracker. This evaluation is conducted on the VOT2018 [14] and VOT2016 [13] datasets, both of which contain annotated object bounding boxes in their video sequences. Higher values in the table indicate more frequent tracker failures after each attack, suggesting a more effective attack performance. Additionally, the average number of failures on the VOT2016 dataset illustrates TrackPGD’s capability in misleading TransT-SEG for object bounding box prediction. The average number of failures is calculated based on the failures on multiple evaluation runs. Despite being primarily designed to generate perturbations based on the network’s mask generation ability, TrackPGD introduces influential perturbations that rank it second in terms of robustness on both the VOT2016 and VOT2018 datasets. RTAA [3], the top-ranked method, is tailored to bounding box degradation using 1024 adversarial bounding boxes, while TrackPGD, aimed at flipping binary mask pixels, still emerges as the second most potent attack for bounding box degradation.
Attack Robustness (VOT2018) Robustness (VOT2016) Ave. failures (VOT2016)
No Attack 1.08 0.54 37.00
IoU [2] 2.14 1.87 115.47
CSA [4] 1.15 0.62 42.53
SPARK [1] 6.60 7.14 283.00
RTAA [3] 24.44 18.63 832.93
TrackPGD 10.76 10.02 483.00
Table 5: The TransT-SEG [11] robustness in predicting object bounding box, attacked by different methods on the VOT2018 [14] and VOT2016 [13] datasets.
We also illustrate the bounding box evaluation of TransT-SEG tracker [11] in several subsequences from VOT2018 [14] in Figure 4. The original annotation is depicted in green, representing a rotated bounding box per frame. The deviation from the actual performance of the TransT-SEG tracker (dark blue box) is comparable to the other attack approaches. The TrackPGD method is indicated in white, which shows a rapid and significant deviation even from the initial frames of the video sequence.
Figure 4: The ground truth is annotated by green and dark blue is the original response of the TransT-SEG [11] tracker. Red, black, yellow, bright blue, and white colors correspond to the SPARK [1], RTAA [3], CSA [4], IoU [2] and TrackPGD attack methods.
5 Conclusion
This paper introduces a novel white-box attack termed TrackPGD, leveraging object binary masks to evaluate the adversarial robustness of new transformer trackers. By employing a binary mask as a new attack proxy, we propose a difference loss comprising two vanilla SegPGD losses to translate a segmentation attack, SegPGD, into a tracking attack, TrackPGD. Our approach addresses two primary challenges: firstly, the binary mask only provides two classes (object and background), unlike the SegPGD method which handles multiple labels in a segmentation map. Secondly, the imbalance between object and background pixels in tracking objects poses a challenge for using just one vanilla SegPGD for attacking transformer trackers. The differential loss enhances attack performance compared to the vanilla SegPGD loss, as demonstrated by the ablation study. Experimental results indicate that TrackPGD achieves superior results in attacking MixFormerM. For OSTrackSTS, our method ranks as the top attack in three metrics – EAO, accuracy, and AUC – out of four measured scores on VOT-STS2022. Furthermore, TrackPGD has also demonstrated excellent performance in altering the bounding box prediction for transformer trackers.
Acknowledgments
This work is supported by the DEEL Project CRDPJ 537462-18 funded by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Consortium for Research and Innovation in Aerospace in Québec (CRIAQ), together with its industrial partners Thales Canada inc, Bell Textron Canada Limited, CAE inc and Bombardier inc.  [^1] https://deel.quebec